OpenStudy (4everaddicted2anime):

Can someone please help me with a geometry question? I will medal and fan

OpenStudy (i_need_help_please):

Maybe. Can you post the question?

OpenStudy (4everaddicted2anime):

Here it is @I_need_help_please Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius.

OpenStudy (i_need_help_please):

Find 1/2 the angle at the center of the right triangle with hypotenuse = 8 and side = 4 --> 30 degrees. Area of the sector = Area of the circle/6 = CA/6 = pi*r^2/6 = 32pi/3 Area of the 2 30-60-90 triangles = 2*bh/2 = 4*8*cos(30) = 16sqrt(3) Subtract the area of the 2 right triangles for CA/6 --> 32pi/3 - 16sqrt(3) [Area of small segment] Large segment area = 64pi - small area = 160pi/3 + 16sqrt(3) sq units

OpenStudy (4everaddicted2anime):

Thank you